Three masses arranged as shown on the right where π‘šπ΄ = 2.00 [kg], π‘šπ΅ = 1.00 [kg], and π‘šπΆ = 3.00 [kg]. What is the location of the center of mass?Β 


1
Expert's answer
2022-05-30T15:18:03-0400

"x=\frac{m_ax_a+m_bx_b+m_cx_c}{m_a+m_b+m_c}=2.33,"

"y=\frac{m_ay_a+m_by_b+m_cy_c}{m_a+m_b+m_c}=3.00."